Scenario-1: A hamburger producing company's variable cost per hamburger is 15.50 taka and fixed cost is 2,20,000/- taka. The company's expected sales of hamburger are 40,000 pieces. Based on this context, answer the following questions: Q1: How many hamburgers the company has to sell to achieve neither profit nor loss situation? (show the step by step calculation) Q2: What is the minimum number of hamburgers the company has to sell to make profit? Q3: How many hamburgers the company has to sell incus minimum amount of loss?


1
Expert's answer
2020-08-28T10:48:58-0400

Question 1


A neither profit nor loss situation is known as a breakeven point.

"Breakeven \space point = \dfrac {Fixed \space Costs}{Contribution \space per \space unit}"


Where,"Unit \space Contribution = selling \space price -unit \space variable \space costs"


However, selling price is not given in the question, therefore we will assume that the firm is making a profit at the planned sales of 40,000 units. We therefore logically expect selling price to be greater than total unit cost at this expected sales level.


"unit \space fixed \space cost = \dfrac {Fixed \space Costs} {number \space of \space units}"

"unit \space fixed \space cost = \dfrac {220,000 \space taka} {40,000 \space pieces}"

"= 5.50 \space taka \space per \space piece"



Total cost per unit = unit variable cost + unit fixed cost

= 15.50 taka + 5.50 taka

= 21 taka per piece.


Therefore, we assume that the selling price per piece exceeds 21 taka. Now, let us assume that the selling price is 35.50 taka per humburger.


Basing on this assumption,

Unit contribution = 35.50 taka - 15.50 taka

= 20 taka per piece


"\therefore \space Breakeven \space point = \dfrac {220,000 \space taka}{20 \space taka}"


"Breakeven \space point = 11,000 \space pieces"


Therefore, assuming a selling price of 35.50 taka per humburger, the company has to sell 11,000 humburger to achieve a neither profit nor loss situation.



Question 2


Since the company achieves neither a profit nor a loss by selling 11,000 humburgers, it means that any additional humburger above 11,000 contributes towards profit. The first 11,000 hamburgers contribute towards covering fixed costs.


To start making a profit, the company has to therefore sell a minimum of 11,001 hamburgers.



Question 3

Since the company fully covers its total costs by selling 11,000 hamburgers, it implies that any sales below 11,000 pieces result in losses. The lower the number of hamburgers sold, the greater the loss as the distance between total costs and total revenue widens. To incur minimum losses, the company, therefore, has to sell only one unit less than the breakeven quantity. Thus, the company has to sell 10,999 pieces.